# Fait: Fault-Aware Fine-Tuning for Better Code Generation

### v1

@@ Abstract.
Modern instruction-tuned large language models (LLMs) have made remarkable progress in code generation. However, these LLMs fine-tuned with standard supervised fine-tuning (SFT) sometimes generate plausible-looking but functionally incorrect code variants. This issue likely stems from the limitation of standard SFT, which treats all tokens equally during optimization and fails to emphasize the error-sensitive segments—specific code differences between correct implementations and similar incorrect variants. To address this problem, we propose Fault-Aware Fine-Tuning (Fait), a novel fine-tuning technique that enhances LLMs’ code generation by (1) extracting multi-granularity (line/token-level) differences between correct and incorrect yet similar implementations to identify error-sensitive segments, and (2) dynamically prioritizing those segments during training via dynamic loss weighting. Through extensive experiments on seven LLMs across three widely-used benchmarks, our method achieves an average relative improvement of 6.9% on pass@1 with just one epoch of training, with some enhanced 6.7B LLMs outperforming closed-source models, e.g., GPT-3.5-Turbo. Furthermore, our fine-tuning technique demonstrates strong generalization with performance improvements ranging from 3.8% to 19.1% across diverse instruction-tuned LLMs, and our ablation studies confirm the contributions of different granularities of differences and loss function components.
1. Introduction
@@
Recently, fine-tuning LLMs using synthetic datasets generated by teacher models has emerged as a popular paradigm for improving code generation capabilities (Luo et al., [n. d.]; Wei et al., 2024; Zheng et al., 2024; Gunasekar et al., 2023; Yu et al., 2024). This paradigm uses stronger teacher models to generate high-quality instruction-response pairs and construct a dataset. These synthetic datasets are then used to fine-tune weaker student models with standard SFT method, which uses instructions to guide LLMs to generate outputs matching reference responses by minimizing cross-entropy loss uniformly across all tokens.
@@
Although these LLMs fine-tuned with standard SFT achieve impressive performance on code generation benchmarks such as HumanEval (Chen et al., 2021), they sometimes generate plausible-looking but incorrect code variants (Huang et al., 2025; Liu et al., 2024a). For example, Llama-3.1-70B-Instruct (Dubey et al., 2024), a model fine-tuned from Llama-3.1-70B using standard SFT, is capable of solving 82.3% of the problems in HumanEval. However, our analysis shows that 34.5% of its failed cases are attributed to deviations in error-prone segments of otherwise correct implementations. As shown in Figure 1, in a vowel identification task, although the LLM successfully implements the core logic of searching for vowels between consonants from the right, it erroneously initializes the starting index of the loop–a mistake in a critical part of the code that renders the program incorrect. We refer to such crucial differences between correct implementations and similar incorrect variants as error-sensitive segments, such as len(word) - 1 and len(word) - 2 in Figure 1. These error-sensitive segments act as critical decision points in code generation, where even slight deviations can determine the correctness of the output.
@@
To address the problem of errors in code generation, we propose Fault-Aware Fine-Tuning (Fait) as a finetuning technique specifically designed to guide LLMs to focus on error-sensitive segments and avoid mistakes in these critical regions, thereby improving the reliability and accuracy of code generation. Implementing this approach involves two main challenges. The first challenge is to identify these error-sensitive segments, as existing instruction-tuning dataset construction methods primarily focus on generating instruction-response pairs without specifically considering these segments (Zhang et al., 2023). To overcome this, we develop a two-phase segment identification component. First, we leverage a teacher model with a carefully designed prompt to generate functionally incorrect yet similar variants of the correct implementations from the existing instruction tuning dataset. Then, we can identify the differences between these correct implementations and their similar yet incorrect variants, which serve as error-sensitive segments. We annotate the tokens in the segments through a multi-granularity method at both line and token levels. Notably, we annotate both the tokens in correct implementations and their corresponding parts in similar yet incorrect variants. The second challenge lies in guiding LLMs to focus on these labeled segments during fine-tuning, as standard SFT treats all tokens with equal importance regardless of their criticality. To address this challenge, we adjust the loss of SFT to prioritize the annotated error-sensitive tokens within correct implementations. Specifically, Fait processes both correct and incorrect implementations to discriminate the error-sensitive segments, and only computes the loss based on correct code implementations. Unlike standard SFT that uniformly weights all tokens during loss computation, we dynamically assign relatively higher weights to those tokens in the correct implementation that correspond to the error-sensitive segments. This methodology enhances LLMs’ capability to discriminate error-sensitive segments when solving programming tasks, thereby increasing the likelihood of generating correct implementation details while suppressing error-prone alternatives.
@@
We validate the effectiveness of the Fait through extensive experiments. Notably, through Fait, the selected LLMs undergo only one epoch of training on their original instruction datasets, yet achieve an average relative improvement of 6.9% on pass@1 (a strict evaluation metric measuring the ratio of first-generated samples that pass all test cases) across three representative code generation benchmarks (Humaneval(+), MBPP(+), and BigCodeBench) (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024b; Zhuo et al., 2024). Among these LLMs, SemCoder-S (Ding et al., 2024) with 6.7B parameters outperforms closed-source models like GPT-3.5-Turbo (OpenAI, 2022) on Humaneval(+) and MBPP(+) benchmarks and MagiCoder-DS with 6.7B parameters outperforms GPT-3.5-Turbo on HumanEval(+). Our method also demonstrates strong generalization capabilities, showing performance improvements ranging from 3.8% to 19.1% across multiple instruction-tuned LLMs, including those trained on closed-source instruction datasets. Moreover, our ablation experiments on Fait confirm the contributions of different granularities of differences in generated code details and loss functions.
@@
We propose a novel framework, Fault-Aware Fine-Tuning (Fait), to effectively guide LLMs to focus on error-prone parts in code. This is achieved by (1) extracting multi-granularity code differences (token-/line-level) to identify error-sensitive segments, and (2) refining SFT to dynamically assign higher weights to these parts during the training process.
@@
2. Approach
Figure 2 illustrates the overview of Fait. This approach takes as input an instruction-tuned LLM and its corresponding instruction-tuning dataset, and outputs an enhanced LLM with improved code generation capabilities that can better discriminate error-sensitive segments. It first augments the dataset by generating similar yet incorrect implementations for each correct response. Then, it identifies error-sensitive segments between the paired implementations and calculates weights for tokens in the correct implementations that differ from the incorrect variants. During fine-tuning, it only computes loss on the correct implementations, with higher weights assigned to tokens in error-sensitive segments, producing an LLM that can better discriminate these error-sensitive segments. The methodology consists of two key components:
@@ 2. Approach
The two components work together to fine-tune LLMs to distinguish between correct implementations and similar yet incorrect alternatives, thereby improving code generation performance. The following subsections detail each component of our methodology.
2.1. Error-Sensitive Segments Identification
The input to this component is the instruction-tuning dataset , where represents the target problem description and denotes the correct implementation. The output is an enhanced dataset with error-sensitive segment information, where represents the corresponding similar but incorrect implementation. To generate incorrect code variants, we utilize a stronger teacher LLM with a carefully designed prompt. The prompt template is shown in Figure 3, which consists of two parts. The first part defines the task for producing incorrect responses corresponding to the target problem and correct answer, explicitly specifying that outputs must be similar to correct solutions, with responses constrained to markdown formatting for consistent post-processing. The second part provides contextual references by presenting the target problem description and the correct solution.
After obtaining pairs of correct and incorrect code, we extract the differences to identify error-sensitive segments and process them at different granularity levels to capture both line-level and token-level information. Specifically, we designate as the pre-change version and as the post-change version.
Line-Level Differences. We align and line-by-line using Python’s difflib library. For each line, the tool assigns a flag indicating whether it should be deleted (-), added (+), or remain unchanged. We extract the lines marked for deletion from and those marked for addition from .
@@ 2.1. Error-Sensitive Segments Identification
Token-Level Differences. We utilize the Levenshtein distance algorithm (Yujian and Bo, 2007) to identify character-level change information between and . The Levenshtein distance algorithm, also known as the edit distance algorithm, quantifies the minimum number of single-character operations (insertions, deletions, or substitutions) required to transform one string into another. We identify the characters that need to be edited to transform the original string () into the modified version (), and record their positions accordingly. For instance, if a character operation is an insertion, we record its position in , as shown in Figure 2. Given that the LLM’s embedding layer is tightly coupled with the LLM’s tokenizer vocabulary, we map these character-level differences to tokens using the LLM’s tokenizer. When character modifications span multiple tokens, all affected tokens are marked.
@@
Hybrid Level Vectors. To create comprehensive representations of error-sensitive segments, we combine line-level and token-level masks. However, these two types of masks operate at different granularities and cannot be directly combined. We must first align these representations to the same granularity to enable their integration.
We convert line-level masks to token-level granularity, as illustrated in Algorithm 1. Lines 1-2 initialize a token-level mask vector (size T) and the position index. Lines 3-16 implement the core logic - for each code line (1 to L), when the line is masked (marked lines, Lines 5-8), all tokens in that line receive mask value 1; otherwise (unmarked lines, Lines 10-15), tokens receive value 0. For and , we identify the corresponding tokens within that line to get and .
With both masks now represented at the token level, we then use an element-wise addition operation to combine and , as follows:
@@
2.2. Dynamic Importance Reweighting
@@
In contrast, the Fait introduces dynamic token-level weights emphasize error-sensitive segments:
@@
Where denotes element-wise multiplication. Inspired by the Bradley–Terry model (Hunter, 2004), a pairwise comparison framework widely used in ranking systems (Baker and Scarf, 2021; Yan, 2016; Menke and Martinez, 2008), we compute dynamic token weights for differentiating tokens in error-sensitive segments:
@@
Where is a hyperparameter controlling the weight range, denotes the mean probability of tokens in error-sensitive segments in , and represents the corresponding value for . This formulation ensures that: (1) When the mean probabilities and are close (indicating the LLM struggles to distinguish the tokens between and ), the weights for differentiating tokens in approach , thereby maximizing emphasis on error-sensitive segments. (2) Conversely, when and diverge significantly (demonstrating the LLM can discriminate the differentiating tokens in and ), the weights diminish toward , reducing emphasis. For tokens shared between and , we assign fixed weights , ensuring the LLM maintains baseline attention to shared elements while prioritizing discriminative features.
@@
This dynamic weighting mechanism guides the LLM to focus on the challenging discriminative aspects of correct implementations, which can improve its code generation capability.
3. Experiments Setup
3.1. Benchmarks
@@
We conduct experiments on three widely used code generation benchmarks to demonstrate the superiority and generality of proposed Fait.
Humaneval (Chen et al., 2021): This benchmark consists of 164 manually crafted programming tasks, created by OpenAI. Each task includes a method signature, a docstring, a method body, and several unit tests. Our work employs both the initial HumanEval and its extended version, HumanEval+(Liu et al., 2024b), which expands the test cases of the original with 80 additional test samples to overcome limitations in test coverage (Liu et al., 2024b).
MBPP (Austin et al., 2021): This benchmark contains 974 Python coding tasks spanning core programming concepts, library utilization capabilities, and more. Our study adopts the extended versions proposed by (Liu et al., 2024b), including MBPP and MBPP+. These collections each contain 378 tasks, with the enhanced version incorporating 35 times the number of test samples.
Bigcodebench (Zhuo et al., 2024): This benchmark presents a rigorous benchmark for code generation, constructed to measure LLMs’ capabilities in utilizing programming tools and the following of complex instructions. It contains 1,140 code-generation problems. In the Complete configuration, each problem provides a function signature, a problem description, and a test suite. A small high-quality subset known as BigCodeBench-Hard contains 148 problems that are more user-centric and challenging. Our study uses both the full set and the hard set, namely, BigCodeBench-Full and BigCodeBench-Hard.
3.2. Metrics
@@
To evaluate code generation performance, we use the Pass@K metric, which is widely used in prior studies (Chen et al., 2021; Mu et al., 2024; Huang et al., 2023). This metric checks whether the generated code passes all test cases successfully within the first K generations. Following prior studies (Fakhoury et al., 2024; Dong et al., 2024; Jiang et al., 2024), our experimental design adopts K=1, focusing exclusively on first-attempt success rates. This metric also aligns with real-world scenarios where developers aim to produce accurate code on the first attempt (Dong et al., 2024). It should be noted that Pass@1 represents a particularly strict evaluation metric for code generation and improving it is challenging. Higher Pass@1 scores indicate better code generation performance.
3.3. Implementation Detail
3.3.1. Data generation
We use Qwen2.5-Coder-32B-Instruct111https://huggingface.co/Qwen/Qwen2.5-Coder-32B-Instruct as the teacher model with temperature=0.8 to generate incorrect code implementations due to its strong coding abilities and good natural language understanding capabilities. To mitigate the potential threat introduced by errors from the teacher model generation, we manually examined a sample of 50 generated outputs. Our analysis shows that the LLM could produce error-sensitive segments as expected: 96% of the generated incorrect samples are similar to the correct ones, while the remaining 4% completely deviate from correct implementations. Due to space limitations, the checked samples are provided in the replication package. This manual inspection helps validate the reliability of our training data and supports the soundness of our experimental findings.
3.3.2. Settings

### v2

# FGit: Fault-Guided Fine-Tuning for Code Generation

@@ Abstract
Modern instruction-tuned large language models (LLMs) have made remarkable progress in code generation. However, these LLMs fine-tuned with standard supervised fine-tuning (SFT) sometimes generate plausible-looking but functionally incorrect code variants. This issue likely stems from the limitation of standard SFT, which treats all tokens equally during optimization and fails to emphasize the error-sensitive segments—specific code differences between correct implementations and similar incorrect variants. To address this problem, we propose Fault-Guided Fine-Tuning (FGit), a novel fine-tuning technique that enhances LLMs’ code generation by (1) extracting multi-granularity (line/token-level) differences between correct and incorrect yet similar implementations to identify error-sensitive segments, and (2) dynamically prioritizing those segments during training via dynamic loss weighting. Through extensive experiments on seven LLMs across three widely-used benchmarks, our method achieves an average relative improvement of 6.9% on pass@1, with some enhanced 6.7B LLMs outperforming closed-source models, e.g., GPT-3.5-Turbo. Furthermore, our fine-tuning technique demonstrates strong generalization with performance improvements ranging from 3.8% to 19.1% across diverse instruction-tuned LLMs, and our ablation studies confirm the contributions of different granularities of differences and hyperparameters.
Index Terms:
LLM, Code Generation, Software EngineeringI Introduction
@@
Recently, fine-tuning LLMs using synthetic datasets generated by teacher models has emerged as a popular paradigm for improving code generation capabilities [1, 2, 3, 4, 5]. This paradigm uses teacher models to generate high-quality instruction-response pairs and construct a dataset. These datasets are then used to fine-tune student models with standard SFT method, which uses instructions to guide LLMs to generate outputs matching reference responses by minimizing cross-entropy loss uniformly across all tokens.
@@
Although these LLMs fine-tuned with standard SFT achieve impressive performance on code generation benchmarks such as HumanEval [6], they sometimes generate plausible-looking but incorrect code variants [7, 8]. For example, Llama-3.1-70B-Instruct [9], a model fine-tuned from Llama-3.1-70B using standard SFT, is capable of solving 82.3% of the problems in HumanEval. However, our analysis shows that 34.5% of its failed cases are largely correct and require modifications in only three or fewer locations to be fixed. We conservatively classify these specific instances as failures attributable to deviations in error-prone segments, highlighting the prevalence of such errors in code generation. As shown in Figure 1, in a task to calculate the product of signs and the sum of magnitudes, although the LLM successfully implements logic similar to the correct version—by correctly handling negative numbers and summing magnitudes-it fails to account for zero. We refer to such crucial differences between correct implementations and similar incorrect variants as error-sensitive segments. These error-sensitive segments act as critical decision points in code generation, where even slight deviations can determine the correctness of the output.
It is crucial to address these errors for code generation. However, standard SFT [10] often overlooks this issue as it does not specifically focus on these segments essential for correctness. Drawing inspiration from curriculum learning [11], which involves training models by progressively moving from easier to more challenging samples to build learning capabilities, we propose Fault-Guided Fine-Tuning (FGit). FGit is a supplementary fine-tuning technique designed to guide instruction-tuned models to further focus their learning on these error-sensitive segments, treating these segments as more challenging samples in a targeted curriculum.
@@
Implementing this approach involves two main challenges. The first challenge is to identify these error-sensitive segments. Systematically collecting large-scale datasets of fine-grained coding errors from real-world scenarios, along with their corresponding correct versions, is difficult. Furthermore, existing instruction-tuning dataset construction methods primarily focus on generating instruction-response pairs without specifically considering these critical segments [12]. Considering traditional code mutation techniques [13, 14] are often constrained by a limited set of predefined transformation rules and may not generate diverse and semantically plausible incorrect variants, we develop a two-phase segment identification component. First, we leverage a teacher model, which can capture a wide range of diverse and semantically plausible error patterns from its extensive code training data, with a carefully designed prompt to generate multiple functionally incorrect yet similar variants of the correct implementations from the existing instruction tuning dataset. From these generated variants, we select the one with the most similarity to the correct implementation and highlight the differences as error-sensitive segments in this paper. We then annotate the tokens in the segments through a multi-granularity method at both line and token levels. Notably, this annotation is applied to both the correct implementations and the selected incorrect variant. The second challenge lies in guiding LLMs to focus on these labeled segments during fine-tuning, as standard SFT treats all tokens with equal importance regardless of their criticality. To address this challenge, we adjust the loss of SFT to prioritize the annotated error-sensitive tokens within correct implementations. Specifically, FGit processes both correct and incorrect implementations to discriminate the error-sensitive segments, and computes loss based on correct code implementations. Unlike standard SFT, which uniformly weights all tokens during loss computation, we dynamically assign relatively higher weights to those tokens in the correct implementation that correspond to the error-sensitive segments. This methodology enhances LLMs’ capability to discriminate error-sensitive segments when solving programming tasks, thereby increasing the likelihood of generating correct implementation details while suppressing error-prone alternatives.
@@
We validate the effectiveness of the FGit through extensive experiments. Notably, through FGit, the selected LLMs achieve an average relative improvement of 6.9% on pass@1 across three representative code generation benchmarks (HumanEval(+), MBPP(+), and BigCodeBench) [6, 15, 16, 17]. Among these LLMs, SemCoder-S [18] with 6.7B parameters outperforms closed-source models like GPT-3.5-Turbo [19] on HumanEval(+) and MBPP(+) benchmarks, and MagiCoder-DS with 6.7B parameters outperforms GPT-3.5-Turbo on HumanEval(+). Our method also demonstrates strong generalization capabilities, showing performance improvements ranging from 3.8% to 19.1% across multiple instruction-tuned LLMs, including those trained on closed-source instruction datasets. Moreover, our ablation experiments on FGit confirm the contributions of different granularities of differences and hyperparameters.
@@
We propose a novel framework, Fault-Guided Fine-Tuning (FGit), to effectively guide LLMs to focus on error-prone parts in code. This is achieved by (1) extracting multi-granularity code differences (token-/line-level) to identify error-sensitive segments, and (2) refining SFT to dynamically assign higher weights to these parts during the training process.
@@
II Approach
Figure 2 illustrates the overview of FGit. This approach takes as input an instruction-tuned LLM and its corresponding instruction-tuning dataset, and outputs an enhanced LLM with improved code generation capabilities that can better discriminate error-sensitive segments. It first augments the dataset by generating similar yet incorrect implementations for each correct response. Then, it identifies error-sensitive segments between the paired implementations and calculates weights for tokens in the correct implementations that differ from the incorrect variants. During fine-tuning, it only computes loss on the correct implementations, with higher weights assigned to tokens in error-sensitive segments, producing an LLM that can better discriminate these error-sensitive segments. The methodology consists of two key components:
1.
@@ II Approach
The two components work together to fine-tune LLMs to distinguish between correct implementations and similar yet incorrect alternatives, thereby improving performance.
II-A Error-Sensitive Segments Identification
The input to this component is the instruction-tuning dataset , where represents the target problem description and denotes the correct implementation. The output is an enhanced dataset with error-sensitive segment information, where represents the corresponding similar but incorrect implementation. To generate incorrect code variants, we utilize a teacher LLM with a carefully designed prompt. We choose an LLM-based approach over code mutation techniques because teacher models, by leveraging error commonalities learned from their extensive code corpora, offer better flexibility in producing a diverse range of incorrect variants. The prompt template is shown in Figure 3, which consists of two parts. The first part defines the task for producing incorrect responses corresponding to the target problem and answer, specifying that outputs should be similar to correct solutions, with responses constrained to markdown formatting for consistent post-processing. The second part provides contextual references to the target problem description and solution.
We generate multiple incorrect variants for each correct implementation and select the one that is most similar to the correct solution. To identify the most similar one, we employ an embedding-based approach because embeddings can capture semantic similarities that purely lexical comparisons might overlook. Specifically, we generate embeddings with UnixCoder [20] for its deep understanding of code structures and semantics derived from its diverse, large-scale code corpora. We then extract the differences to identify error-sensitive segments and process them at different granularity levels to capture both line-level and token-level information. Specifically, we designate as the pre-change version and as the post-change version.
Line-Level Differences. We align and line-by-line using Python’s difflib library. For each line, it assigns a flag indicating whether it should be deleted (-), added (+), or remain unchanged. We extract the lines marked for deletion from and those marked for addition from .
@@ II-A Error-Sensitive Segments Identification
Token-Level Differences. We utilize the Levenshtein distance algorithm [21] to identify character-level change information between and . The Levenshtein distance algorithm, also known as the edit distance algorithm, quantifies the minimum number of single-character operations (insertions, deletions, or substitutions) required to transform one string into another. We identify the characters that need to be edited to transform the original string () into the modified version (), and record their positions accordingly. For instance, if a character operation is an insertion, we record its position in , as shown in Figure 2. Given that the LLM’s embedding layer is tightly coupled with the LLM’s tokenizer vocabulary, we map these character-level differences to tokens using the LLM’s tokenizer. When character modifications span multiple tokens, all affected tokens are marked.
@@
Hybrid Level Vectors. To create comprehensive representations of error-sensitive segments, we combine line-level and token-level masks. However, these two types of masks operate at different granularities and cannot be directly combined. To get line-level masks with token-level granularity, we first initialize a new token-level mask vector with all zeros, corresponding to the total number of tokens in the code. Then, for each line in the original code, all tokens belonging to that line are assigned the line’s mask value (1 if the line is marked, 0 if it is unmarked) in this new token-level mask. This process yields the token-level representations and when applied to and respectively. We then use an element-wise addition operation to combine and , as follows:
@@
II-B Dynamic Importance Reweighting
@@
In contrast, FGit introduces dynamic token-level weights emphasize error-sensitive segments:
@@
Where denotes element-wise multiplication. Inspired by the Bradley–Terry model [22], a pairwise comparison framework widely used in ranking systems [23, 24, 25], we compute dynamic token weights for differentiating tokens in error-sensitive segments:
@@
where is a hyperparameter controlling the weight range, denotes the mean probability of tokens in error-sensitive segments in , and represents the corresponding value for . This formulation ensures that: (1) When the mean probabilities and are close (indicating the LLM struggles to distinguish the tokens between and ), the weights for differentiating tokens in approach , thereby maximizing emphasis on error-sensitive segments. (2) Conversely, when and diverge significantly (demonstrating the LLM can discriminate the differentiating tokens in and ), the weights diminish toward , reducing emphasis. For tokens shared between and , we assign fixed weights , ensuring the LLM maintains baseline attention to shared elements while prioritizing discriminative features. To be noted that this dynamic weight is differentiable, as and are derived from the model’s output probabilities for and within the error-sensitive segments. Consequently, through gradient updates, the optimization process reinforces , also implicitly steering the model away from generating .
@@
This dynamic weighting mechanism guides the LLM to focus on challenging discriminative aspects of correct implementations, which can improve its code generation capability.
III Experiments Setup
III-A Benchmarks and Metrics
@@
We conduct experiments on three widely used code generation benchmarks to demonstrate the superiority and generality of FGit. We use HumanEval(+), MBPP(+) [16, 6, 15] and Bigcodebench[17]. Specifically, our study uses both the full set and hard set with complete configuration of Bigcodebench, namely, BigCodeBench-Full and BigCodeBench-Hard.
@@
To evaluate performance, we use the Pass@K metric, which is widely used in prior studies [6, 26, 27]. Following prior studies [28, 29, 30], our experimental design adopts K=1, focusing exclusively on first-attempt success rates. This metric also aligns with real-world scenarios where developers aim to produce accurate code on the first attempt [29].
III-B Implementation Detail
III-B1 Data generation
We use Qwen2.5-Coder-32B-Instruct [31] as the teacher model with temperature=0.8 to generate three incorrect code implementations for each sample. Our rationale for selecting this model is threefold. First, generating an incorrect solution by referencing a correct one is comparatively less demanding on a model’s capabilities than generating a correct solution from scratch. Secondly, our task benefits from a model with strong coding abilities, and Qwen2.5-Coder-32B-Instruct possesses high coding proficiency; for instance, its performance on HumanEval(+) and MBPP(+) even surpasses that of GPT-4o. Lastly, using open-source models also facilitates reproducibility.
To investigate whether the teacher model can generate implementations that are incorrect and similar to the correct solutions, we employ both manual inspection and automated quantitative validation. First, two authors independently examine a sample of 50 generated outputs, which are randomly sampled from the Evol-instruct dataset, assessing them based on two criteria: functional correctness and similarity to the original correct code. A discussion is held to resolve the disagreements. We do not invite others because all the disagreements are resolved during discussion. The Cohen’s Kappa coefficient [32] is 0.78. Our analysis shows that the LLM could produce similar yet incorrect samples as expected: 94% of the generated samples are similar to the correct ones yet incorrect, while the remaining 6% introduce changes that do not affect correctness. To quantitatively analyze the correctness and similarity of the generated responses, we select benchmarks with test cases for further validation. Specifically, we choose HumanEval and MBPP to generate incorrect yet similar responses and use Unixcoder[20] to calculate cosine similarity. We find that 94.1% of generated responses are incorrect using the test cases in benchmarks, and the average similarity score is 0.95, aligning with our needs. Due to space limitations, the manually checked samples and generated responses are in our replication package.
III-B2 Settings